Ques. 1 The Sunrise enterprises, a single product company, provides you the following data for the Month of Sep. 2020.

·       Sales (5,000 units @ $20/unit): $100,000

·       Contribution Margin per unit: $12

·       Total fixed expenses for the month: $18,000


Required: Calculate break-even point( in units and in dollars) and margin of safety( in Percentage, units and Dollars) for Sunrise enterprises using above data. Also draw a CVP graph and show the sales volume representing break-even point and margin of safety on the graph.




1
Expert's answer
2020-11-12T17:35:32-0500

Break-even point in units is:

"BEP = \frac{18,000} {20 - 12} = 2,250" units.

In dollars is:

"BEP = \frac{18,000} {(20 - 12)/20} = 45,000" dollars.

Margin of safety is: "\frac{100,000 - 45,000} {100,000} × 100 = 55" % or $55,000.